ochemistry Find Ksp for Tl2S using E° for TI and TI2S (the latter is -0.964 V) 1.
Answer #1

Using the E=E - 0.0591 log ksp squation, i Th₂ s n 2 + +82- [m=2 Ezolfor equillibrium) 0= -0.964-0.059 1 log kop 0.964x2 = lo